# Morpho-Photometric Classification of KiDS DR5 Sources Based on Neural Networks: A Comprehensive Star-Quasar-Galaxy Catalog

### v1

@@ 4.1.1 Internal Testing Dataset
Although the validation dataset was not directly involved in model training, its role in model selection may inadvertently lead to overfitting, thereby increasing the risk of information leakage. To address this concern and objectively assess the performance of our model, we randomly selected 20,000 spectroscopically confirmed sources from the SDSS-KiDS cross-matched sample to compose an independent testing dataset. The sample was exclusively for final testing, without participating in the training or validation stages. This ensures an unbiased assessment, mitigating the potential for the model to overfit to any particular subset of the data.
@@
On this isolated testing dataset, our multimodal neural network achieved an overall accuracy of 98.76%, closely aligning with the 98.69% accuracy observed in the validation dataset. Such consistency indicates that our model generalizes well from the validation phase to unseen new data, which is crucial for practical applications. We then conducted more detailed performance assessments on each class of sources, including Precision, Recall, and F1-score. The results are presented in Figure 3 and Table 1.
@@
The classification metrics revealed that galaxies exhibited the best performance with all three metrics exceeding 99%. This success is largely attributed to the enormous galaxy sample size used during training. This was followed by the stars, with all metrics also exceeding 98.6%, reflecting robust identification of these two types of celestial objects. QSOs achieved slightly lower Precision and Recall at 96.71% and 94.53% respectively, likely due to the relatively smaller training sample size, as well as the potential for misclassification of some low-redshift active galactic nuclei as galaxies due to contamination from their host galaxies. This latter point is evidenced by the confusion matrix depicted in Figure 4, where the majority of misclassified QSOs were grouped with galaxies. Despite this, a robust F1-score of 95.6% demonstrates that the model is still reliable in terms of QSO classification.
@@
In machine learning paradigms, our model assigns a probability vector to each target for classification, with each value representing the likelihood of a predefined class (such as Stars, QSOs, or Galaxies in our definition). The sum of these probabilities is 1, and typically, a higher probability (p) for a target being classified into a certain class indicates greater confidence, which implies higher Purity (Precision). Therefore, we can set a threshold for $p$, retaining only those samples that exceed this threshold as the final result, thereby enhancing the sample Purity. It is important to note that this strategy, while beneficial for sample Purity, may decrease the Completeness of the sample. For example, a higher threshold may exclude objects with insufficient classification confidence, potentially leading to the absence of some actual category members in the final catalog.
@@
To balance Purity and Completeness in various astronomical applications, we tested the performance of our multimodal neural network at different thresholds. Figure 3 shows the variations in evaluation metrics as a function of the threshold, with some detailed values listed in Table 1. As expected, an increase in the threshold can improve Precision, Recall, and the F1-score, whereas $\rm Frac_{thre}$ and the derived Completeness ($\rm Frac_{thre}\times$ Recall, as shown in Table 1) exhibit a declining trend. We also examined the confusion matrix at different thresholds (Figure 5), and the results showed a clear decrement in misclassification rates with threshold increments. This trend suggests that raising the threshold for $p$ indeed amplifies the Purity of the sample but at the cost of Completeness.
@@
An analysis of Figures 3 and 5 reveals that the choice of probability threshold significantly impacts the classification of QSOs, in contrast to a minor effect on the classification of Galaxies. This is primarily because the model already achieves excellent performance in predicting Galaxies, leaving limited scope for further refinement. Intriguingly, we found that when the threshold is fixed at 0.9, the Completeness for all categories can be maintained above 90%, and the Purity exceeds 98%. As the threshold is further increased to 0.99, Purity can reach above 99%, while Completeness still maintains a level of 82%. Beyond this point, Completeness quickly declines despite a marginal gain in Purity. Hence, setting the probability threshold between 0.9 to 0.99 appears to offer an effective balance between Purity and Completeness.
Figure 6: Redshift distribution for the GAMA testing dataset. The black solid line shows the distribution for all targets, while the blue dotted line, red dashed line, and green solid line represent the distributions for targets predicted as stars, quasars, and galaxies by the multimodal neural network, respectively.
Figure 7: $gri$ color-composited images of some misclassified targets. The vertical axis represents the spectral classification labels from SDSS, and the horizontal axis represents the predictions made by the multimodal neural network.
Figure 8: Color-color diagrams ($g-r$ vs. $r-i$) for KiDS and SDSS datasets. From top to bottom, the panels represent stars, QSOs, and galaxies. The red shaded regions indicate the cross-matched data between SDSS and KiDS, while the blue contours represent all KiDS data brighter than 23 mag.
@@ 4.1.2 External Testing Datasets
Considering the issue of pre-selected sources in the SDSS dataset, there is potential for our neural network model to overfit to this data. To comprehensively evaluate the generalization capability of the model in practical applications, it is necessary to validate it on external datasets. Currently, most publicly released spectroscopic survey data have limited overlap with KiDS, which constrains the availability of extensive testing samples. We primarily utilize Gaia and GAMA as external data sources to verify the generalization performance of our model on stellar and galaxy (including QSO) classification, respectively.
@@
Most targets in the Gaia database do not have corresponding spectroscopic observations, but its high-precision astrometric measurements enable us to filter out a pure stellar sample. This sample is considered unbiased in terms of stellar radiation and can effectively avoid potential selection effects. Through cross-matching with KiDS, we obtained over 3.4 million independent testing samples (see Section 2.3). None of these data were involved in model tuning or evaluation. The test results showed that as high as 99.74% of the targets were correctly predicted as stars, even exceeding the performance on the internal testing dataset. This discrepancy is mainly attributed to the contamination present in SDSS spectroscopic classifications (see Section 4.2), indicating that our algorithm actually has superior performance in star classification.
@@
For extragalactic objects, we adopted the abundant spectroscopic redshift galaxy sample from GAMA for testing. Over 230k of these targets that overlap with KiDS but not with SDSS. The prediction results revealed that only 97.70% of the targets were categorized as galaxies, with the remaining 1.21% classified as stars and 1.11% as quasars, which did not appear to meet expectations. However, upon further examination of the redshift distribution of these targets (see Figure 6), we found that the majority of those predicted as stars had redshifts around 0, confirming that they indeed reside within the Milky Way. After removing these low-redshift (arbitrarily chosen $z<$ 0.04) targets, the fraction of categorized as galaxies increased to 98.72%, while stars and quasars accounted for 0.26% and 1.02%, respectively. Figure 6 also shows that targets with redshifts greater than 0.8 were almost all classified as QSOs. Given the depth of the GAMA survey, it is unlikely to observe such high-redshift galaxies, so these classification results should be reasonable. If QSO classification is also regarded as correct, we can conclude that the classification Precision for extragalactic objects reached 99.74%.
In this test, we did not impose any restrictions on the threshold of $p$, but the model still demonstrated excellent performance for both Galactic and extragalactic targets, further confirming the stability and reliability of our algorithm across different datasets. This establishes a solid foundation for future applications to more extensive samples and other survey data.
4.2 Limitations and Challenges
In addition to the choice of ML algorithms, data quality fundamentally determines the upper limit of model performance. Clean data allows the model to focus on learning the underlying patterns that are consistent with the real-world problems during the training process, thereby enhancing its stability and interpretability. Conversely, dirty data may mislead the learning process of model parameters and evaluation results, increasing the risk of overfitting. Indeed, there are some misclassified targets in the SDSS spectral data (Bolton et al., 2012; Lyke et al., 2020). However, since the proportion of these misclassifications is relatively low (about only a few percent), the model still performs impressively well on the testing dataset. This can also be attributed to the strong noise-resistant capability of DNNs, as well as the use of techniques such as regularization and batch training during the training process. Through random inspection of misclassified samples, we found that some SDSS labels themselves were erroneous, and our multimodal neural network model was even able to effectively correct these errors, especially for galaxies misclassified as stars by SDSS (as shown in Figure 7).
The performance of the model is also influenced by the data volume. A larger dataset can provide richer and more detailed feature representations, which helps the model learn complex patterns and effectively mitigate noise and outlier. This, in turn, enhances its accuracy and generalization ability. In our training, the number of quasar samples was significantly less than that of stars and galaxies, which somewhat limited the recognition capabilities of the model for these targets. This indicates that although our method has demonstrated impressive classification performance, there is still room for further improvement as spectral data continues to accumulate in the future.
@@ 4.2 Limitations and Challenges
To determine the applicable brightness range of the model, we conducted the following tests. First, we selected two equal-sized samples based on their $r$-band magnitudes, with brightness ranges of $18.516.5\leq r<$ 18.5 mag and $20.518.5\leq r<$ 20.5 mag, respectively. By comparing the model performance of these two brightness groups, we found that their classification abilities were basically equivalent. Even when extrapolating the testing data by 1 mag toward the faint end, the performance evaluation metrics did not show a significant decline for either model. Subsequently, we further tested the model using three different brightness thresholds ($r<19$ mag, $r<20$ mag, and $r<21$ mag), confirming that all three results could essentially extrapolate by 1 mag under these conditions. These two tests help us infer that our multimodal neural network can effectively handle targets with magnitudes up to $\sim$23 mag.
@@ 6 $p_{\rm max}$ Maximum output probability - double
Note. — Description of columns in the KiDS DR5 Galaxy-QSO-Star Catalog. The table lists the properties stored for each object, including identifiers, coordinates, magnitudes, classification results, and probabilities.
@@ 4.3 Application to KiDS DR5
We apply the trained multimodal neural network model to the entire KiDS DR5 dataset to predict the categories of target sources. Our goal is to compile a new Star-QSO-Galaxy catalog covering an area of approximately 1350 deg2. According to our prior testing outcomes, we only classify targets with an $r$-band magnitude less than 23. To facilitate diverse research interests, we provide the $r$-band magnitude and the associated maximum output probability ($p_{\rm max}$) for each target, with detailed information listed in Table 2.
This value-added catalog contains a total of 27,334,751 targets, of which 6,427,859 are identified as stars, 3,304,636 as quasars, and 17,602,220 as galaxies. Notably, 13,068,232 targets fall within the $22<r\leqslant 23$ mag range in the $r$-band, which are extrapolated results, accounting for nearly half ($\sim$47.81%) of the total number of targets. This portion of the data exceeds the brightness range directly trained by our model, yet the model has demonstrated acceptable accuracy and stability within this magnitude range in previous tests. Therefore, we decide to include these extrapolated results in the final catalog.
To validate the effectiveness of our final classification results, we compared the $g-r$ vs. $r-i$ color-color distributions of stars, QSOs, and galaxies in our KiDS catalog with those in the SDSS dataset. As shown in Figure 8, the results exhibit a general agreement between the two datasets, suggesting that our model successfully captures the overall color properties of different sources. Furthermore, we examined the size-magnitude diagram for each class of sourcesfor each class of sources, which is a commonly used method to differentiate between stars and galaxies (e.g., Zuntz et al., 2018). Here, we adopt the half-light radius ($R_{1/2}$) to represent size, and the magnitude is in the $r$-band. The results are shown in Figure 9. It is evident that stars are primarily located at the bright end, with sizes typically consistent with the PSF, reflecting their point-source nature. In contrast, galaxies are predominantly found at the fainter end and exhibit significantly larger sizes compared to stars, indicative of their more extended structures. Quasars are primarily distributed between the star and galaxy populations in the size-magnitude space. This distribution aligns well with known observational characteristics and supports our classification results.
@@
Figure 11: Examples of spectra, SEDs, and images for different categories. The left panels show the SDSS spectra (black solid line) and the SEDs from KiDS nine bands (red points). The right panels show color-composited images from KiDS $gri$-bands.

### v2

@@ 4.1.1 Internal Testing Dataset
Although the validation dataset was not directly involved in model training, its role in model selection may inadvertently lead to overfitting, thereby increasing the risk of information leakage. To address this concern and objectively assess the performance of our model, we randomly selected 20,000 spectroscopically confirmed sources from the SDSS-KiDS cross-matched sample to compose an independent testing dataset. The sample was exclusively for final testing, without participating in the training or validation stages. This approach provides a more objective assessment by mitigating biases introduced during the training process and model selection.
@@
On this isolated testing dataset, our MNN achieved an overall accuracy of 98.76%, closely aligning with the 98.69% accuracy observed in the validation dataset. Such consistency indicates that our model generalizes well from the validation phase to unseen new data, which is crucial for practical applications. We then conducted more detailed performance assessments on each class of sources, including Precision, Recall, and F1-score. The results are presented in Figure 3 and Table 1.
@@
The classification metrics revealed that galaxies exhibited the best performance with all three metrics exceeding 99%. This success can be attributed to the large training sample size and the distinct morphological features of galaxies. This was followed by the stars, with all metrics also exceeding 98.6%, reflecting robust identification of these two types of celestial objects. QSOs achieved slightly lower Precision and Recall at 96.71% and 94.53% respectively, likely due to the relatively smaller training sample size, as well as the potential for misclassification of some low-redshift active galactic nuclei (AGNs) as galaxies due to contamination from their host galaxies. This latter point is evidenced by the confusion matrix depicted in Figure 4, where the majority of misclassified QSOs were grouped with galaxies. Despite this, a robust F1-score of 95.6% demonstrates that the model is still reliable in terms of QSO classification.
@@
In ML paradigms, our model assigns a probability vector to each target for classification, with each value representing the likelihood of a predefined class (such as Stars, QSOs, or Galaxies in our definition). The sum of these probabilities is 1, and typically, a higher probability (p) for a target being classified into a certain class indicates greater confidence, which implies higher Purity (Precision). Therefore, we can set a threshold for $p$, retaining only those samples that exceed this threshold as the final result, thereby enhancing the sample Purity. It is important to note that this strategy, while beneficial for sample Purity, may decrease the Completeness of the sample. For example, a higher threshold may exclude objects with insufficient classification confidence, potentially leading to the absence of some actual category members in the final catalog.
@@
To balance Purity and Completeness in various astronomical applications, we tested the performance of our MNN at different thresholds. Figure 3 shows the variations in evaluation metrics as a function of the threshold. As expected, an increase in the threshold can improve Precision, Recall, and the F1-score, whereas $\rm Frac_{thre}$ and the derived Completeness ($\rm Frac_{thre}\times$ Recall) exhibit a declining trend. This trend suggests that raising the threshold for $p$ indeed amplifies the Purity of the sample but at the cost of Completeness.
@@
An analysis of Figures 3 reveal that the choice of probability threshold significantly impacts the classification of QSOs, in contrast to a minor effect on the classification of Galaxies. This is primarily because the model already achieves excellent performance in predicting Galaxies, leaving limited scope for further refinement. Intriguingly, we found that when the threshold is fixed at 0.9, the Completeness for all categories can be maintained above 90%, and the Purity exceeds 98%. As the threshold is further increased to 0.99, Purity can reach above 99%, while Completeness still maintains a level of 82%. Beyond this point, Completeness quickly declines despite a marginal gain in Purity. Hence, setting the probability threshold between 0.9 to 0.99 appears to offer an effective balance between Purity and Completeness.
@@ 4.1.2 External Testing Datasets
Considering the issue of pre-selected sources in the SDSS dataset, there is potential for our neural network model to overfit to this data. To comprehensively evaluate the generalization capability of the model in practical applications, it is necessary to validate it on external datasets. Currently, most publicly released spectroscopic survey data have limited overlap with KiDS, which constrains the availability of extensive testing samples. We primarily utilize Gaia and GAMA as external data sources to verify the generalization performance of our model on stellar and galaxy classification, respectively. For QSOs, while GAMA provides partial coverage of high-redshift samples, a uniformly selected and sufficiently large pure QSO dataset is currently lacking. Therefore, instead of relying on a comprehensive spectroscopic validation dataset, we assess the effectiveness of the model in QSO identification by comparing the number of QSO candidates it recovers across the survey footprint.
@@
Most targets in the Gaia database do not have corresponding spectroscopic observations, but its high-precision astrometric measurements enable us to filter out a pure stellar sample. This sample is considered unbiased in terms of stellar radiation and can effectively avoid potential selection effects. Through cross-matching with KiDS, we obtained over 3.4 million independent testing samples (see Section 2.3). None of these data were involved in model tuning or evaluation. The test results showed that as high as 99.74% of the targets were correctly predicted as stars, even exceeding the performance on the internal testing dataset. This discrepancy may be attributed to the contamination present in SDSS spectroscopic classifications (see Section 5.2) and the fact that GAIA dataset includes brighter sources, which could hint at the superior performance of our algorithm in star classification under these conditions.
Figure 6: Redshift distribution for the GAMA testing dataset. The black solid line shows the distribution for all targets, while the blue dotted line, red dashed line, and green solid line represent the distributions for targets predicted as stars, quasars, and galaxies by the MNN, respectively.
@@
For galaxies, we adopted the abundant spectroscopic redshift galaxy sample from GAMA for testing. Over 230k of these targets that overlap with KiDS but not with SDSS. The prediction results revealed that only 97.70% of the targets were categorized as galaxies, with the remaining 1.21% classified as stars and 1.11% as quasars, which did not appear to meet expectations. However, upon further examination of the redshift distribution of these targets (see Figure 6), we found that the majority of those predicted as stars had redshifts around 0, confirming that they indeed reside within the Milky Way. After removing these low-redshift (arbitrarily chosen $z<$ 0.04) targets, the fraction of categorized as galaxies increased to 98.72%, while stars and quasars accounted for 0.26% and 1.02%, respectively. Figure 6 also shows that targets with redshifts greater than 0.8 were almost all classified as QSOs. Given the depth of the GAMA survey, it is unlikely to observe such high-redshift galaxies, so these classification results should be reasonable. If QSO classification is also regarded as correct, we can conclude that the classification Recall for extragalactic objects reached 99.74%.
To estimate the number of QSOs expected in the KiDS footprint under a given apparent magnitude limit, we adopt the method described in Section 2.2 to select QSOs in the Stripe 82 region of the SDSS [^4] We use the subregion defined by $32^{\circ}<{\rm R.A.}<45^{\circ}$ and $-1\fdg 25<{\rm Decl.}<1\fdg 25$, where the QSO surface density is highest. We select all spectroscopically confirmed QSOs in this region (excluding duplicate observations), and compute the QSO surface density as a function of apparent magnitude. To obtain a robust estimate, we randomly sample 100 independent 4 deg2 subregions and adopt the median surface density as the final value., where spectroscopic completeness is relatively high (Yang & Shen, 2023). At $r<21$, the QSO surface density derived from SDSS is $\sim$83 deg-2, corresponding to $\sim$112.5k QSOs over the 1350 deg2 KiDS footprint. Our MNN model predicts a comparable number of $\sim$119.5k QSO candidates, indicating good agreement. At $r<22$, the SDSS data yield an estimated $\sim$209.5k QSOs, while our model identifies a significantly higher number of $\sim$394.5k—nearly a factor of two larger. This discrepancy likely reflects the rapid decline in SDSS spectroscopic completeness at the faint end.
[…]
In the above test, we did not impose any restrictions on the threshold of $p$, but the model still demonstrated excellent performance for both Galactic and extragalactic targets. This result supports the stability and reliability of our algorithm, at least for bright-end targets. Nakoneczny et al. (2021) employed the XGBoost method to construct a reliable QSO sample from KiDS DR4, limited to sources with $r<22$. We cross-matched their catalog with our predictions and found that 79% of their QSOs were also classified as QSOs by our model. Most of the mismatched objects have $p<0.9$, and after removing them, the disagreement rate drops to $\sim$6%. These findings establish a solid foundation for future applications to more extensive samples and other survey data.
Table 2: Classification Performance at Different Magnitude Ranges.
Train Range Test Range Star QSO Galaxy
Precision Recall F1-score Precision Recall F1-score Precision Recall F1-score
$r<21$ $21<r\leqslant 22$ 0.9237 0.9685 0.9456 0.9552 0.8723 0.9118 0.9806 0.9910 0.9858
$r<20$ $20<r\leqslant 21$ 0.9647 0.9799 0.9722 0.9064 0.9177 0.9120 0.9851 0.9792 0.9822
$r<19$ $19<r\leqslant 20$ 0.9634 0.9820 0.9726 0.9685 0.9036 0.9349 0.9684 0.9869 0.9776
$r<18$ $18<r\leqslant 19$ 0.9838 0.9861 0.9849 0.9711 0.8837 0.9253 0.9835 0.9949 0.9892
$19<r\leqslant 21$ $21<r\leqslant 22$ 0.9231 0.9638 0.9430 0.9444 0.8797 0.9109 0.9822 0.9895 0.9858
$18<r\leqslant 20$ $20<r\leqslant 21$ 0.9659 0.9609 0.9634 0.8814 0.9207 0.9006 0.9848 0.9767 0.9807
$17.5<r\leqslant 19.5$ $20.5<r\leqslant 20.5$ 0.9638 0.9820 0.9728 0.9472 0.9305 0.9387 0.9803 0.9807 0.9805
$17<r\leqslant 19$ $19<r\leqslant 20$ 0.9231 0.9638 0.9430 0.9444 0.8797 0.9109 0.9822 0.9895 0.9858
Note. — Evaluation metrics (Precision, Recall, and F1-score) for stars, quasars, and galaxies trained on datasets with different $r$-band magnitude ranges.
4.2 Testing the Extrapolation Performance of the Model
@@
To evaluate the applicable brightness range of the model, we divided the cross-matched SDSS data into multiple subsets based on their $r$-band magnitudes and trained the model on each subset to assess its extrapolation performance. The training (and validation) datasets were constructed using two different approaches:
  1. 1.
    Threshold-based division: Four brightness thresholds were defined, and for each threshold, all sources brighter than the given value were grouped into a single training dataset.
  2. 2.
    Bin-based division: The data were divided into four bins with equal magnitude intervals, and all sources within each bin were used as an independent training dataset.
For each trained model, we evaluated its performance using datasets fainter than the corresponding training dataset by a series of thresholds ranging from 0.5 to 1.5 mag. The results, summarized in Table 2, indicate that the model maintains acceptable performance when extrapolating up to 1.0 mag beyond the training range. For example, the F1-scores for stars, QSOs, and galaxies consistently exceed 94%, 90%, and 97%, respectively, across all tests. This suggests that our MNN has the potential to handle targets with magnitudes up to $\sim$23 mag. However, additional testing is required to fully characterize its extrapolation limitations. Therefore, sources with $r$-band magnitudes brighter than 22 mag are regarded as reliable targets, while those with magnitudes between 22 and 23 mag are treated as acceptable extrapolated targets.
Figure 7: Color-color diagrams ($g-r$ vs. $r-i$) for KiDS and SDSS datasets. From top to bottom, the panels represent stars, QSOs, and galaxies. The left panels show objects with $r\leqslant 22$, while the right panels correspond to objects with $22<r\leqslant 23$. The red shaded regions indicate the cross-matched data between SDSS and KiDS, and the green contours represent all KiDS data within the corresponding magnitude range.
Figure 10: Examples of spectra, SEDs, and images for different categories. The left panels show the SDSS spectra (black solid line) and the SEDs from KiDS nine bands (red points). The SEDs agree well with the spectra for star and QSO, while galaxy photometric points tend to be higher due to the larger KiDS aperture compared to the SDSS fiber. The right panels show color-composited images from KiDS $gri$-bands.
@@ 6 $p_{\rm max}$ Maximum output probability - double
Note. — Description of columns in the KiDS DR5 Galaxy-QSO-Star Catalog. The table lists the properties stored for each object, including identifiers, coordinates, magnitudes, classification results, and probabilities. (The full catalog can be accessed at https://cosviewer.com/datasets/kids-dr5-target-classify. The $r$-band magnitudes will be updated after the photometry data from KiDS DR5 is fully released.)
@@ 4.3 Application to KiDS DR5
We apply the trained MNN model to the entire KiDS DR5 dataset to predict the categories of target sources. Our goal is to compile a new Star-QSO-Galaxy catalog covering an area of approximately 1350 deg2. According to our prior testing outcomes, we only classify targets with an $r$-band magnitude less than 23. To facilitate diverse research interests, we provide the $r$-band magnitude and the associated $p_{\rm max}$ for each target, with detailed information listed in Table 3.
This value-added catalog contains a total of 27,335,836 targets, of which 6,549,300 are identified as stars, 2,804,654 as quasars, and 17,981,882 as galaxies. Notably, 13,068,547 targets fall within the $22<r\leqslant 23$ mag range in the $r$-band, which are extrapolated results, accounting for nearly half ($\sim$47.81%) of the total number of targets. This portion of the data exceeds the brightness range directly trained by our model, yet the model has demonstrated acceptable accuracy and stability within this magnitude range in previous tests. Therefore, we decide to include these extrapolated results in the final catalog.
To validate the effectiveness of our final classification results, we compared the $g-r$ vs. $r-i$ color-color distributions of stars, QSOs, and galaxies in our KiDS catalog with those in the SDSS dataset. As shown in Figure 7, the reliable targets exhibit good agreement with the SDSS dataset, whereas the extrapolated targets display some deviations but remain largely consistent. This demonstrates that our model effectively captures the overall color properties of different sources. Furthermore, we examined the size-magnitude diagram for each class of sources for each class of sources, which is a commonly used method to differentiate between stars and galaxies (e.g., Zuntz et al., 2018). Here, we adopt the half-light radius ($R_{1/2}$) to represent size, and the magnitude is in the $r$-band. The results are shown in Figure 8. It is evident that stars are primarily located at the bright end, with sizes typically consistent with the PSF, reflecting their point-source nature. In contrast, galaxies are predominantly found at the fainter end and exhibit significantly larger sizes compared to stars, indicative of their more extended structures. Quasars are primarily distributed between the star and galaxy populations in the size-magnitude space. This distribution aligns well with known observational characteristics and supports our classification results.
@@
5 Discussion